Does there pass a plane through the lines (x+4)/3 =y/2 =(z−1)/3
and x/2 =(y−1)/1 =(z+1)/1?Justify.
1
Expert's answer
2020-12-17T19:39:30-0500

Two lines in space (R3) can be in three states: parallels, intersect, skew.

If they are skew lines, they do not lie in the same plane.

Check which type they belong to:

1)parallel

Direction vector (L1=(x+4)/3 =y/2=(z−1)/3) is of the form n=(3,2,3)

Direction vector (L2=x/2 =(y−1)/1 =(z+1)/1) is of the form m=(2,1,1)

n it is not equal to a*m (a∈R) So they are not parallel.

2)intersect

if these two lines intersect then they have a common point A=(x1, y1, z1)

this point satisfies both equations of lines

(x1+4)/3 = y1/2 =( z1−1)/3

x1/2 =( y1−1)/1 =( z1+1)/1


Find the coordinates of the point using a system of two equations

1)x1/2 =( y1−1)/1

2)(x1+4)/3 = y1/2

1) x1=2*y1-2

2)(2*y1-2+4)/3 = y1/2

(2*y1+2)/3 = y1/2

4*y1+4=3*y1

y1=-4

x1=2*(-4)-2=-10

substitute these coordinates in both equations and find for each z1

1)-4/2=( z1−1)/3

2)-10/2=z1+1

1)z1=-6+1=-5

2)z1=-5-1=-6

The solutions don't match so the lines don't intersect.

So they skew and it follows that the plane cannot pass through them.